NatalieBM:

Marcella ate one cup of air-popped popcorn. Calories 92.8 Fat 1.1 g Carbohydrates 18.7 g Sodium 2.2 mg Which equation can Marcella use to find how many carbohydrates she consumed? c = 18.7 divided by 3 c = 3 times 18.7 c = 3 divided by 18.7 c = 1.1 times 3

supie:

Since she ate one cup meaning \(\frac{1}{3}cup\) of air popped popcorn.

supie:

For every whole air popped popcorn there is 18.7g Carbohydrates Meaning that \(\frac{18.7g\ of\ Carbs}{popcorn}\) (When I say popcorn I mean the whole bag [but Marcella is eating one cup of the whole thing]) So, \(18.7g\ of\ carbs\ per(/)\ popcorn\)

supie:

\(\color{aqua}{\text{Originally Posted by}}\) @supie Since she ate one cup meaning \(\frac{1}{3}cup\) of air popped popcorn. \(\color{aqua}{\text{End of Quote}}\) Back to this, It says \(\frac{1}{3}\) meaning that since it says `Carbohydrates` `18.7g` and fractions is division it would be \(c=18.7 ? ? Notice that fraction mean division and the number at the bottom is what we will be dividing it by. That should show you what goes in the two `? ?` spots.